OpenStudy (anonymous):

how do you solve this? 9-5z=12-(6z=7) what is z equal to

jigglypuff314 (jigglypuff314):

ummm do you mean 9-5z=12-(6z+7)?

jigglypuff314 (jigglypuff314):

9 - 5z = 12 - (6z + 7) distribute the negative so 9 - 5z = 12 - 6z - 7 9 - 5z = -6z + 5 then add 6z to both sides +6z +6z 9 + z = 0 + 5 then subtract both sides by 9 -9 -9 z = 5 - 9 so z =